# Applications and resource estimates for open system simulation on a quantum computer

### v1

@@ II.3 MagLab single material \$ cost
LANL’s MagLab pipeline contains costly experiments with high magnetic fields. The cost of that pipeline is known (\$39M a year is the budget of MagLab), and it is also known that there is a quantum-classical separation in a Ca3Co2O6 experiment [nekrashevich2022reaching] proposed as a benchmark. The open system contribution to the dynamics of that benchmark is likely to be important for reproducing the experimental behavior and eventually replacing the experiment altogether.
@@
The specific experiment concerning the nonequilibrium spin states does not have direct applications. In general, spin materials are studied for the advancement of:
@@ II.4 Metal-Insulator transition material search \$ cost
In this section, we discuss the utility of a material search for the new materials with Metal-Insulator Transition (MIT). First of all, we note that this transition happens at room temperature, so it may be possible to predict it ab initio classically, but existing attempts were not very successful. Our utility estimate will be valid either way, but it does not have to be quantum computers that succeed in this material search. A table of possible applications of MIT can be found in the supplement of [lee2024nanoelectronics], including:
@@
  • Artificial Synapse
  • Artificial Neuron
  • Photodetector
  • Gas Sensor
@@
While most research focuses on making a transistor out of MIT material, it is not expected that it would replace conventional semiconductor technology for integrated circuits. Indeed, in the reviews discussing continuing Moore’s law, MIT materials are not listed as a potential candidate for a new transistor for conventional computers. There is a field, however, where it is considered promising: the field of neuromorphic computing. Indeed, the reviews [tang2019bridging] would list it as the 3rd best out of 6 potential candidate technologies, giving the weight factor of 1/5. There is no direct estimate for the market size of neuromorphic computing, but we could use AI market share of semiconducting hardware listed in [batra2019artificial] as 20%. The US semiconductor market is \$80b, while the global one is \$600b. We do not expect that the chip production using a completely new paradigm of neuromorphic computing can easily be reverse-engineered over a 10-year timescale, assuming that the details of the fabrication process are kept private early on, which means we can use the global revenue for our estimate. As the main financial investment in semiconductor manufacturing is the expensive fabrication facilities, not the research into the transistor technology itself, we estimate a 1/10 ratio of how the chip technology will factor in the revenue. Finally, we need to estimate the probability of switching from the current dedicated AI hardware to neuromorphic computing. Again, since there are various competing technologies (GPU, TPU, etc.), we estimate another 1/10 ratio as the neuromorphic devices take the last place out of 4 candidates. The resulting utility estimate is:
$\text{Utility}=\frac{1}{5}\times\frac{1}{5}\times\frac{1}{10}\times\frac{1}{10}\times\$600\text{b}=\$24\text{M}$ (3)
@@
The size of the research grants spent on this search so far can be estimated from the reports on DOE programs and European collaborations. Q-MEEN-C in San Diego has been funded for eight years with \$3M per year, focusing on MIT materials and spintronics for neuromorphic computing. It also contains experimental effort in fabricating device prototypes. We expect roughly 1/10 of funding to go to the computational efforts, resulting in a \$2.4M scientific funding estimate of utility. We also note the Materials Project, from which we learn that about 100 materials with MIT are known [georgescu2021database], though the funding structure of that project does not allow an easy estimate, and 1.5M euro FIRSTORM program as an example of a smaller scale computational research program on this topic. It is clear that many other computational research programs on this topic were funded, raising the total scientific funding utility to 10x the amount we found (an order of magnitude estimate), \$20M. We note that the two estimates have produced a comparable number.
@@ II.5 Gibbs state preparation: most likely targets
In this section, we discuss a subroutine of many quantum algorithms: using a thermal state as an initial state. Applications that allow for those initial states are the material searches such as the high-Tc and MIT material searches, as well as quantum chemistry problems that involve modeling materials such as surface chemistry. Some problems in fundamental science also require thermal initial states, such as the MagLab experiment on Ca3Co2O6 described above. We estimate the total utility of all the applications where Gibbs state prep is a reliable tool to be of the order of \$100M and refer to companion publications for specific resource and utility estimates. While Gibbs state prep alone does not solve those applications, and each individual task in those \$100M can be only \$2M or less, with varying required number of runs, it may well be possible that Gibbs state preparation will dominate the resource cost. The reasons why one would still choose Gibbs state preparation instead of other approaches are explained below.
@@
We first digress to discuss a more common problem of ground state preparation. To obtain the desired precision, existing algorithms of ground state preparation usually contain an unknown factor in the dependence of the total resource cost on the system size. For example, a ground state estimation algorithm can enhance the probability of the ground state for an initial state with nonzero overlap, with circuit size growing with the inverse overlap. The initial state is then constructed by another circuit, and the cost of that circuit is impossible to compute before the quantum computer actually becomes available. It is sometimes assumed that this circuit’s cost is negligible, so only the ground state estimation resource cost is counted in the resource estimate. It is, in general, not known what is the class of Hamiltonians where a polynomial accuracy can be obtained in polynomial time. 1D matrix product states (MPS) of gapped Hamiltonians are one example of such a system [dalzell2019locally]: it is proven that the MPS can be computed classically and then constructed as a circuit inside the quantum computer. The overheads of such a construction may be quite big, though. Besides the rigorously proven results, we also have intuition about the behavior of the quantum annealing approach to the preparation of the polynomial overlap state. Essentially, a path avoiding first-order phase transitions needs to be found connecting the target state to some accessible initial state, such as the lower overlap state prepared by other methods or the product state. First-order phase transitions, however, are quite common. Thus, the adiabatic state preparation cannot be applied blindly. Every target system’s phase diagram needs to be analyzed independently. We see that there is no unified method for providing a resource estimate for the state preparation of application-relevant states. The best one can hope for is an extrapolation for small system sizes, where both positive and negative results have been claimed [lee2023evaluating] since the accuracy of the extrapolation is not controlled by anything.
@@
In comparison, thermal state preparation allows for more accessible guarantees. Indeed, the thermal state of any system can be prepared for sufficiently high temperatures [alhambra2021locally]. Moreover, a bound exists [chen2021fast] for the time to prepare the thermal state for all systems satisfying the Eigenstate Thermalization Hypothesis (ETH). Extrapolating the parameters in that bound from the exact diagonalization results would be more stable than extrapolating the overlaps in ground state estimation since those parameters reflect physical phenomena. The distinction between ETH and non-ETH systems is also well understood, with the latter containing many-body localization, spin glasses, and integrable models. Most of the application models are not in any of those categories. For example, glassy effects have never been observed in nonequilibrium electron dynamics of large molecules in quantum chemistry. That suggests that we could reliably prepare finite temperature states in those systems by a universal algorithm. Compare it with the adiabatic ground state preparation, which relied on avoiding first-order phase transitions above. To the best of our knowledge, this is the only reliable way to obtain resource counts for the initial state preparation algorithms.
@@ III Problem Statement
by specifying the Hamiltonian $H$, initial state $\rho(0)$ and the Lindblad generators $L_{\mu}$. While in our application instances, we will use simple local forms of $L_{\mu}$, a variation of the same application instances that has a correct thermal steady state [chen2023quantum] and is more likely to produce valuable results uses $L_{\mu}^{\text{therm}}$ defined via Hamiltonian evolution $e^{iHt}$ of a simple local operator. Such an operator would generally span the entire system and contain exponentially many Pauli’s, yet it can be evaluated efficiently on a quantum computer [chen2023quantum]. It is this operator that is used in our third, broad-scope application instance of preparing Gibbs states, which we do not discuss further as it directly follows [chen2023quantum].
@@ III.1 MagLab’s application problem (Ca3Co2O6)
Quantum-Classical separation can be argued for from the results of [nekrashevich2022reaching]. In particular, in Fig. 7, the results of QMC are shown as a proxy for nonequilibrium open system dynamics. Our goal is to reproduce it. The corresponding experiment is shown in Fig. 1.
@@
We note that while many MagLab studies correspond to closed system processes, the specific experiment shown in Fig. 1 is a nonequilibrium open system evolution. The model is a version of the Transverse Field Ising model. The longitudinal and transverse magnetic fields are controlled by the experimentalist. The spins are arranged in the following three sublattices as shown in Fig. 2.
@@
$J_{1}=-1$ is ferromagnetic, while $J_{2}=0.1|J_{1}|$ and $J_{3}=0.1|J_{1}|$ are antiferromagnetic. The overall energy scale is $23.9(2)$K, which can be translated into $4.98\cdot 10^{11}$z. The z-component of the Hamiltonian is given by:
@@
An additional term is the transverse field $H_{x}=h_{TF}\sum_{i}\sigma_{i}^{x}$ with $h_{TF}=|J_{1}|/300$. The total Hamiltonian is $H(t)=H_{z}(t)+H_{x}$. The problem is to simulate the nonequilibrium states arising during the hysteresis loop: a sweep of the longitudinal field $h(t)$. The longitudinal field $h(t)$ is swept from 00 to $1.4|J_{1}|$ over a week of real time, $T=5.57\cdot 10^{5}$s $=2.77\cdot 10^{17}$ in the units where $|J_{1}|=1$ in the original experiment. We will only consider $T=1$s$=4.97\cdot 10^{11}$ in natural units as it will turn out to be enough to illustrate how long nonequilibrium dynamics is still out of reach of quantum simulation. The schedule of sweeps is shown in the figure above.
@@
The nonequilibrium values of the z-magnetization arise as the result of imperfect thermal relaxation. The only classical method that reproduces them is quantum Monte Carlo, which theoretically shouldn’t have predictive power for nonequilibrium dynamics. Sometimes, it still works, but it cannot by itself be conclusive. A quantum simulator is the only tool that would potentially lessen MagLab’s workload and also answer the question it cannot access.
Here, we specify the open system addition to the above Hamiltonian that would reproduce the data as shown. The simplest way that may already describe the data is to add Lindblad operators $\sim\sigma^{-}=|0\rangle\langle 1|$ and $\sigma^{+}=|1\rangle\langle 0|$ conditioned on the states of neighboring spins, with rates given by the classical energies of the computational basis spin configurations.
@@
and $P_{j,s_{j}}=\frac{1}{2}(\sigma^{z}_{j}+s_{j})$ is a projector onto a state $|s_{j}\rangle$. The energies $H_{z}^{(i)}(s,s_{0})$ are the terms of $H_{z}$ containing $\sigma^{z}_{i}$, after the substitution $\sigma^{z}_{j}\to s_{j}$ for all neighbors $j\in NN(i)$ in the interaction graph, and $\sigma^{z}_{i}\to s_{0}$. The inverse temperature $\beta$ is taken from the schedule in Fig. 1. Here we joined the spin flips for particular configurations of the neighbors that happen to have equal rate into a single $L_{\mu}$ operator, so that it would preserve the coherence between different states of the neighbors in that case. $P_{r_{\pm}}$ are the corresponding joined projectors that may have rank higher than $1$. $r_{+}$ and $r_{-}$ iterate through all possible values of the rates $r_{\pm,s}$. We note that $r_{\pm}=1$ is the most frequent as well as the highest rate, with about a half of all configurations. The number of terms $L_{i,1,+}$ and $L_{i,1,-}$ is the range of index $i$, which is the total number of spins.
Due to a very small value of $h_{TF}$, this choice of $L_{\mu}$ may be sufficient to reproduce the nonequilibrium plateau observed in the experiment, and either positive or negative result will shed light on the underlying physics of the phenomenon. A more advanced approach that is more likely to reproduce the experiment is using $L_{\mu}^{\text{therm}}$ as defined in [chen2023quantum].

### v2

@@ II.3 MagLab single material \$ cost
MagLab pipeline contains costly experiments with high magnetic fields. The cost of that pipeline is known (\$39M a year is the budget of MagLab), and it is also known that classical algorithms do not fully explain Ca3Co2O6 experiment [nekrashevich2022reaching] proposed as a benchmark. The open system contribution to the dynamics of that benchmark is likely to be important for reproducing the experimental behavior and eventually replacing the experiment altogether.
@@
The specific experiment [nekrashevich2022reaching] concerning the nonequilibrium spin states does not have direct applications. In general, spin materials are studied for the advancement of:
@@ II.4 Metal-Insulator transition material search \$ cost
In this section, we discuss the utility of a material search for new materials with Metal-Insulator Transition (MIT). First of all, we note that this transition happens at room temperature, so it may be possible to predict it ab initio classically, but existing attempts were not very successful. Our utility estimate will be valid either way, but it does not have to be a quantum computer that succeeds in this material search. A table of possible applications of MIT can be found in the supplement of [lee2024nanoelectronics], including:
@@
  • Artificial Synapses
  • Artificial Neurons
  • Photodetectors
  • Gas Sensors
@@
While most research focuses on making a transistor out of MIT material, it is not expected that it would replace conventional semiconductor technology for integrated circuits. Indeed, a literature search in the reviews discussing the continuation of Moore’s law showed that MIT materials are not listed as potential candidates for a new transistor for conventional computers. There is a field, however, where it is considered promising: the field of neuromorphic computing. Indeed, the reviews [tang2019bridging] would list it as the 3rd best out of 6 potential candidate technologies, giving the weight factor of 1/5. There is no direct estimate for the market size of neuromorphic computing, but we could use the entire AI market share, which is 20% of the semiconducting hardware market as listed in [batra2019artificial]. The US semiconductor market is \$80b, while the global one is \$600b. Ten years from now, it is projected to grow to \$1,100b. We do not expect that the chip production using a completely new paradigm of neuromorphic computing can easily be reverse-engineered over a 10-year timescale, assuming that the details of the fabrication process are kept private early on, which means we can use the global revenue for our estimate. As the main financial investment in semiconductor manufacturing is the expensive fabrication facilities, not the research into the transistor technology itself, we estimate a 1/10 ratio of how the chip technology will factor in the revenue. Finally, we need to estimate the probability of switching from the current dedicated AI hardware to neuromorphic computing. Again, since there are various competing technologies (GPU, TPU, etc.), we estimate another 1/10 ratio as the neuromorphic devices take the last place out of 4 candidates. The resulting utility estimate is:
$\text{Utility}=\frac{1}{5}\times\frac{1}{5}\times\frac{1}{10}\times\frac{1}{10}\times\$1,100\text{b}=\$440\text{M}$ (3)
@@
The size of the research grants spent on this search so far can be estimated from the reports on DOE programs and European collaborations. Q-MEEN-C in San Diego has been funded for eight years with \$3M per year, focusing on MIT materials and spintronics for neuromorphic computing. It also contains experimental effort in fabricating device prototypes. We expect roughly 1/10 of funding to go to the computational efforts, resulting in a \$2.4M scientific funding estimate of utility. We also note the Materials Project, from which we learn that about 100 materials with MIT are known [georgescu2021database], though the funding structure of that project does not allow an easy estimate, and 1.5M euro FIRSTORM program as an example of a smaller scale computational research program on this topic. An automated estimate of the total size of funded computational research programs on this topic can be performed following the methodology of [agrawal2024quantifyingfaulttolerantsimulation]. We leave it for future work, assuming the result would come out to \$40M in the programs funded today. As only 100 materials were found, and for the full search, we assume that 1000 materials need to be studied, we expect that the total cost would match the $\sim$\$400M industry estimate.
@@ II.5 Gibbs state preparation: most likely targets
In this section, we discuss a subroutine of many quantum algorithms: using a thermal state as an initial state. Applications that allow for those initial states are the material searches such as the high-Tc and MIT material searches, as well as quantum chemistry problems that involve modeling materials such as surface chemistry. Some problems in fundamental science also require thermal initial states, such as the MagLab experiment on Ca3Co2O6 described above. We estimate the total utility of all the applications where Gibbs state preparation is a reliable tool to be of the order of \$100M and refer to companion publications [watts2024fullereneencapsulatedcyclicozonegeneration, otten2024quantumresourcesrequiredbinding, nguyen2024quantumcomputingcorrosionresistantmaterials, bellonzi2024feasibilityacceleratinghomogeneouscatalyst, penuel2024feasibilityacceleratingincompressiblecomputational, elenewski2024prospectsnmrspectralprediction, agrawal2024quantifyingfaulttolerantsimulation, bartschi2024potentialapplicationsquantumcomputing, saadatmand2024faulttolerantresourceestimationusing] for specific resource and utility estimates. While Gibbs state preparation alone does not solve those applications, and each individual task in those \$100M can be only \$2M or less, with varying required number of runs, it may well be possible that Gibbs state preparation will dominate the resource cost. The reasons why one would still choose Gibbs state preparation instead of other approaches are explained below.
@@
We first digress to discuss a more common problem of ground state preparation. To obtain the desired precision, existing algorithms of ground state preparation usually contain an unknown factor in the dependence of the total resource cost on the system size. For example, a ground state estimation algorithm can enhance the probability of the ground state for an initial state with nonzero overlap, with circuit size growing with the inverse overlap. The initial state is then constructed by another circuit, and the cost of that circuit is impossible to compute before the quantum computer actually becomes available. It is sometimes assumed that this circuit’s cost is negligible, so only the ground state estimation resource cost is counted in the resource estimate. It is, in general, not known what is the class of Hamiltonians where the ground state can be obtained to polynomial accuracy in polynomial time. 1D matrix product states (MPS) of gapped Hamiltonians are one example of such a system [dalzell2019locally]: it is proven that the MPS can be computed classically and then constructed as a circuit inside the quantum computer. The overheads of such a construction may be quite big, though. Besides the rigorously proven results, we also have intuition about the behavior of the quantum annealing approach to the preparation of the polynomial overlap state. Essentially, a path avoiding first-order phase transitions needs to be found connecting the target state to some accessible initial state, such as the lower overlap state prepared by other methods or the product state. First-order phase transitions, however, are quite common. Thus, the adiabatic state preparation cannot be applied blindly. Every target system’s phase diagram needs to be analyzed independently. We see that there is no unified method for providing a resource estimate for the state preparation of application-relevant states. The best one can hope for is an extrapolation for small system sizes, where both positive and negative results have been claimed [lee2023evaluating] since the accuracy of the extrapolation is not controlled by anything.
@@
In comparison, thermal state preparation allows for more accessible guarantees. Indeed, the thermal state of any system can be prepared for sufficiently high temperatures [alhambra2021locally]. Moreover, a bound exists [chen2021fast] for the time to prepare the thermal state for all systems satisfying the Eigenstate Thermalization Hypothesis (ETH). Extrapolating the parameters in that bound from the exact diagonalization results would be more stable than extrapolating the overlaps in ground state estimation since those parameters reflect physical phenomena. The distinction between ETH and non-ETH systems is also well understood, with the latter containing many-body localization, spin glasses, and integrable models. Most of the application models are not in any of those categories. For example, glassy effects have never been observed in nonequilibrium electron dynamics of large molecules in quantum chemistry. That suggests that we could reliably prepare finite temperature states in those systems by a universal algorithm. Compare it with the adiabatic ground state preparation, which relied on avoiding first-order phase transitions above. To the best of our knowledge, thermal state preparation is the only reliable way to obtain resource counts for the initial state preparation algorithms.
@@ III Problem Statement
by specifying the Hamiltonian $H$, initial state $\rho(0)$ and the Lindblad generators $L_{\mu}$. While most of our estimates use simple local forms of $L_{\mu}$, only a carefully picked $L_{\mu}$ has a correct thermal steady state. For instance, [chen_quantum_2023] uses $L_{\mu}^{\text{therm}}$ defined via Hamiltonian evolution $e^{iHt}$ of a simple local operator. While such an operator would generally span the entire system and contain exponentially many Pauli’s, it can be evaluated in polynomial time on a quantum computer. The prefactors of that approach, however, can be prohibitive, as we will show for the average case of MIT material search. We believe that a more practical approach would be to optimize $L_{\mu}$ on a classical computer such that it produces a thermal state approximately while maintaining a simple local form. In the third, broad-scope application instance of preparing Gibbs states, we expect dedicated Gibbs state preparation algorithms (also proposed in [chen_quantum_2023], as well as [cubitt2023dissipative, chen2021fast]) to be more efficient than a simulation of an open system evolution. In all of these algorithms, the algorithm runtime necessary for equilibration is a property of the chosen approach and is not guaranteed to be polynomial except in special cases of high temperature or low dimensionality. In principle, it can be extrapolated from the sizes that can be addressed classically. If those sizes prove too small, one can use the formula in [chen2021fast] relating the relaxation time to the parameters of the eigenstate thermalization hypothesis, which may be determined at smaller sizes. We have not investigated the necessary resource estimates further in this work.
@@ III.1 MagLab’s application problem (Ca3Co2O6)
We choose this experiment as a target for quantum simulation because Quantum-Classical separation can be argued for from the results of [nekrashevich2022reaching]. Specifically, in Fig. 6 of that work, the results of Quantum Monte-Carlo (QMC) are shown as a proxy for experimental Fig. 7, showing nonequilibrium open system dynamics. Our goal is to reproduce Fig. 7 of [nekrashevich2022reaching] via an ab initio simulation. That is in contrast to QMC, which is a classical random process that only qualitatively reproduces the quantum nonequilibrium open system dynamics. In particular, the relationship between the number of spin updates and the time elapsed in the quantum dynamics, as well as the discretization of the imaginary time register, are fitting parameters when QMC is used as a proxy for nonequilibrium dynamics. They cannot be computed ab initio.
@@
For the purposes of this experiment, Ca3Co2O6 is well described by a version of the Transverse Field Ising model. The temperature and the longitudinal and transverse magnetic fields are controlled by the experimentalist, as shown in Fig. 1. We stress that while many MagLab studies correspond to closed system processes, the specific experiment shown in Fig. 1 is a nonequilibrium open system evolution.
@@
The spins are arranged in three sublattices, as shown in Fig. 2. $J_{1}=-1$ is ferromagnetic, while $J_{2}=0.1|J_{1}|$ and $J_{3}=0.1|J_{1}|$ are antiferromagnetic. The overall energy scale is $23.9(2)$K, which can be translated into $4.98\cdot 10^{11}$z. The z-component of the Hamiltonian is given by:
@@
An additional term is the transverse field $H_{x}=h_{TF}\sum_{i}\sigma_{i}^{x}$ with $h_{TF}=|J_{1}|/300$. The total Hamiltonian is $H(t)=H_{z}(t)+H_{x}$. The problem is to simulate the nonequilibrium states arising during the sweep of the longitudinal field $h(t)$. The longitudinal field $h(t)$ is swept from 00 to $1.4|J_{1}|$ over a week of real time, $T=5.57\cdot 10^{5}$s $=2.77\cdot 10^{17}$ in the units where $|J_{1}|=1$ in the original experiment. We will only consider $T=1$s$=4.97\cdot 10^{11}$ in natural units as it will turn out to be enough to illustrate how long nonequilibrium dynamics is still out of reach of quantum simulation.
@@
Here, we specify the open system addition to the above Hamiltonian that aims to reproduce the data shown in Fig. 7 of [nekrashevich2022reaching]. The simplest way that may already describe the data is to add Lindblad operators $\sim\sigma^{-}=|0\rangle\langle 1|$ and $\sigma^{+}=|1\rangle\langle 0|$ conditioned on the states of neighboring spins, with rates given by the classical energies of the computational basis spin configurations.
@@
and $P_{j,s_{j}}=\frac{1}{2}(\sigma^{z}_{j}+s_{j})$ is a projector onto a state $|s_{j}\rangle$. The energies $H_{z}^{(i)}(s,s_{0})$ are the terms of $H_{z}$ containing $\sigma^{z}_{i}$, after the substitution $\sigma^{z}_{j}\to s_{j}$ for all neighbors $j\in NN(i)$ in the interaction graph, and $\sigma^{z}_{i}\to s_{0}$. The inverse temperature $\beta$ is taken from the schedule in Fig. 1. Here, we joined the spin flips for particular configurations of the neighbors that happen to have equal rates into a single $L_{\mu}$ operator so that it would preserve the coherence between different states of the neighbors in that case. $P_{r_{\pm}}$ are the corresponding joined projectors that may have a rank higher than $1$. $r_{+}$ and $r_{-}$ iterate through all possible values of the rates $r_{\pm,s}$. We note that $r_{\pm}=1$ is the most frequent as well as the highest rate, with about half of all configurations. The number of terms $L_{i,1,+}$ and $L_{i,1,-}$ is the range of index $i$, which is the total number of spins. The number of distinct Lindblad terms on each spin can be estimated as follows. From Fig. 2 we see that there are 2 neighbors linked by $J_{1}$ and 12 neighbors linked by $J_{2}=J_{3}$. There are $m+1$ possible values of the total spin for $m$ spins. Thus, the number of possible rates can be estimated as $3\cdot 13$ (we took into account that half of them lower the energy and thus map to $1$).
Due to a very small value of $h_{TF}$, this choice of $L_{\mu}$ may be sufficient to reproduce the nonequilibrium states observed in the experiment [nekrashevich2022reaching], and either positive or negative result will shed light on the underlying physics of the phenomenon. A more advanced approach that is more likely to reproduce the experiment is using $L_{\mu}^{\text{therm}}$ as defined in [chen_quantum_2023].